OpenStudy (anonymous):

Evaluate the Integral:

OpenStudy (goformit100):

Wherre is it ?

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 2x^4-4x^3+13x^2-6x+10 }{ (x-2)(x^2-2x+5) }dx\]

OpenStudy (anonymous):

I just need help with the partial fraction decomposition.

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

i am doing hw for pfd also so i may be able to work on this with you

OpenStudy (anonymous):

TI-89

OpenStudy (anonymous):

TI-89 cant be used in my class lol. let me try and figure this out @diddo525 i think that since this is a high degree in the neumirator we will have to do division first

OpenStudy (anonymous):

No it's not. Expand the bottom. The denominator is greater than the numerator for the degree.

OpenStudy (anonymous):

well its only \[x^3-4x^2+9x-10\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Never mind. I got it. It's really really ugly though >.< .

OpenStudy (anonymous):

man i hate to say it but im stumped, we just started learning it, could you explain?

OpenStudy (anonymous):

Thanks anyways :) .

OpenStudy (anonymous):

Allright. The denominator has an irreducible quadratic right?

OpenStudy (anonymous):

Because we can't factor it in any way.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So forget about that for now. We can rewrite the fraction as: \[\frac{ A }{x-2 } + \frac{ Bx+c }{ x^2-2x+5 }+\frac{ Dx+E }{ (x^2-2x+5)^2 }\]

OpenStudy (anonymous):

Whenever we have an irreducible quadratic, we can rewrite the numerator with a linear function like I did.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

@sjerman1

OpenStudy (anonymous):

sorry yes, yes it does

OpenStudy (anonymous):

\[(-20A-10B+9C-2D+E)x+(14A+9B-4c+D)x^2+(-4A-4B+c)x^3+(A+B)x^4+25A-10C-2E\]

OpenStudy (anonymous):

We set each term equal to the corresponding degree in the original polynomial: